### v1

@@ Abstract
Decentralized prediction markets (DePMs) allow open participation in event-based wagering without fully relying on centralized intermediaries. We review the history of DePMs which date back to 2011 and includes hundreds of proposals. Perhaps surprising, modern DePMs like Polymarket deviate materially from earlier designs like Truthcoin and Augur v1. We use our review to present a modular workflow comprising seven stages: underlying infrastructure, market topic, share structure and pricing, trading, market resolution, settlement, and archiving. For each module, we enumerate the design variants, analyzing trade-offs around decentralization, expressiveness, and manipulation resistance. We also identify open problems for researchers interested in this ecosystem.
@@ 1 Introduction
Noticing these points of differences inspired us to ask what are all the design decisions involved in creating a DePM? We reviewed over a hundred projects, distilled into the 35+ notable DePMs listed in Appendix 0.A. From these, we break the DePM design space into a ‘modular workflow’ (the method used in an SoK at AFT 2021 on oracles [24]) with seven stages: underlying infrastructure, market topic, share structure and pricing, trading, market resolution, settlement, and archiving. For each stage, we enumerate the possible designs and discuss competing trade-offs.
@@ 2 Preliminaries
2.1 Methodology
We obtained a collection of academic works on decentralized prediction markets, as well as various intersecting topics including (centralized) prediction markets, oracles, DeFi, and AMMs. We used our knowledge of the field, Google Scholar (search and cited by features), and citations within papers. Our library is available, sorted by topic, on Zotero. [^2] Zotero: https://www.zotero.org/groups/5750510/2024_prediction_markets We also identified projects without white-papers or academic papers, reviewing the 97 current projects listed on a recently released online dashboard [67] (noting that some listed are DePM-adjacent rather than true DePMs) and by studying 20+ historical systems that are no longer active. We also searched news sources for opinions and issues on leading decentralized prediction markets, such as Polymarket. We used affinity diagrams to cluster design decisions into the seven stages of our final modular framework. Relevant artifacts and data are available on GitHub. [^3] GitHub: https://github.com/MadibaGroup/2024-PredictionMarkets
2.2 Taxonomy
@@
A wager is a two-party contract with payouts based on the outcome of a future event. Consider Alice and Bob who wager on the same outcome of an event. With a fixed-odds bookmaker (or online betting), Alice’s contract is different from Bob’s contract in at least two regards: (i) it specifically names Alice as the counterparty and (ii) the payouts could be different if the odds changed between Alice’s wager and Bob’s. By contrast, in a prediction market contract (called a outcome share), Alice and Bob hold identical contracts: (i) all contracts are between the market operator and whoever redeems the contract, and (ii) the payout is exactly the same (typically \$0 if incorrect and \$1 if correct). Odds are reflected in the price paid for a prediction market contract (i.e., variable cost and fixed payout), while a bookmaker contract has a fixed cost and variable payout. Thus the key distinction is that prediction market outcome shares are fungible and can be freely traded between participants, enabling a free market that communicates information to the public through outcome share prices, trading volume, market depth, and other financial market metrics. Other wager structures include parimutuel betting [72, 44, 74], back/lay exchanges [17] and cash for difference contracts [23].
@@ 2.3 Definitions
A (single) market is a tuple $M=(E,\Omega,J,R)$, where $E$ is a well-defined uncertain event, $\Omega$ is a nonempty outcome space for $E$, $J$ is a finite index set of contract labels (“shares”), and $R=(R_{j})_{j\in J}$ are nonnegative payoff functions with $R_{j}:\Omega\to\mathbb{R}_{\geq 0}$. When $M$ resolves to $\omega_{M}\in\Omega$, one unit of share $j\in J$ pays $R_{j}(\omega_{M})$ units of $\mathcal{N}$ (see Def. 2).
@@
For a market $M=(E,\Omega,J,R)$, suppose there exists a bijection $\iota:J\to\Omega$ and $R_{j}(\omega)\in\{0,1\}$ with $\sum_{j\in J}R_{j}(\omega)=1$ for all $\omega\in\Omega$. Then $M$ is a winner-take-all (WTA) or Arrow–Debreu market: a unit claim of label $j$ pays $1$ iff the realized outcome equals $\iota(j)$, and 0 otherwise.
@@
A prediction–market system is a tuple $\mathcal{S}=(\mathcal{M},\mathcal{N},\mathsf{Res})$, where $\mathcal{M}$ is a countable set of markets, $\mathcal{N}$ is a numeraire (unit of account), and $\mathsf{Res}=\{\mathrm{res}_{M}\}_{M\in\mathcal{M}}$ is a family of resolution registers such that, for each $M$ with outcome space $\Omega_{M}$, we have $\mathrm{res}_{M}\in\{\bot\}\cup\Omega_{M}$, $\mathrm{res}_{M}$ is initially $\bot$, and $\mathrm{res}_{M}$ transitions exactly once to some $\omega_{M}\in\Omega_{M}$.
@@
For every market $M=(E,\Omega,J,R)\in\mathcal{M}$ operating in system $\mathcal{S}=(\mathcal{M},\mathcal{N},\mathsf{Res})$, the following axioms hold.
@@
    Issuance and Solvency. Let $S_{j}(M)\geq 0$ be the outstanding supply of label $j\in J$, and let $\mathsf{Treas}_{M}\geq 0$ be the market’s treasury (in the numeraire $\mathcal{N}$). The system maintains, at all times,
    $\mathsf{Treas}_{M}\;\geq\;\sup_{\omega\in\Omega_{M}}\;\sum_{j\in J}S_{j}(M)\,R_{j}(\omega).$ (1)
    Thus when $q\geq 0$ new shares of any $j\in J$ are issued, treasury is increased sufficiently:
    $\Delta\mathsf{Treas}_{M}\;\geq\;\sup_{\omega\in\Omega_{M}}\;q\,R_{j}(\omega).$ (2)
@@
    Transfer and Fungibility. Holdings of each label $j\in J$ are transferable between accounts and transfers conserve per–label totals $S_{j}(M)$. For each $j\in J$, shares are indistinguishable: for any $\omega\in\Omega_{M}$ and $q\geq 0$, redeeming $q$ units yields $q\,R_{j}(\omega)$.
@@
    Settlement and redemption. The resolution register satisfies $\mathrm{res}_{M}\in\{\bot\}\cup\Omega$, is initially $\bot$, and transitions exactly once [^4] Real world DePMs like Polymarket might resolve a market, receive a dispute of over the outcome, and resolve it differently after a process (see Section 3.5). In the definition, resolution refers to the final outcome only. An outcome is final when shares can be redeemed for payouts. to a realized outcome $\omega_{M}\in\Omega$. Afterward, any holder of $q\geq 0$ units of label $j\in J$ may redeem them for $q\,R_{j}(\omega_{M})$ units of $\mathcal{N}$. Redemption decreases outstanding shares $S_{j}(M)^{\prime}\leftarrow S_{j}(M)-q$ and debits the treasury $\mathsf{Treas}_{M}^{\prime}\leftarrow\mathsf{Treas}_{M}-q\,R_{j}(\omega_{M})$, preserving the solvency invariant above.
@@ 2.4 An Example of a Market
Also of interest is how the prediction market did not obviously extract insider information which is in violation of what theory would predict [34]. The director did state he did not participate in the market and advised his team working on the film not to either [8]. Friction for novice users is also high: web3 apps have a learning curve and if insiders were based at HBO in the US, access would require circumvention of Polymarket’s geofencing. Perhaps these reasons kept insiders out of the market for the 5 days it ran.
@@ 3 Modular Workflow
Underlying Infrastructure §3.1Market Topic §3.2Share Structure and Pricing §3.3Trading §3.4Market Resolution §3.5Settlement §3.6Archiving §3.7
We now turn to the design landscape of DePMs and step through our modular workflow (cf. [24]), summarized above. Some design decisions will be common issues for both centralized and decentralized prediction markets. We include these anyways for completeness. However we put the emphasis on discussing design decisions that are pertinent to decentralization and permissionlessness.
@@ 3.2 Market Topic
Table 2: Some pitfalls that illustrate the difficulty in properly defining a prediction market topic.
@@
Careful attention must be paid to both the general topic of the market and the ‘fine-print’ or exact predicate that decides the market. Table 2 provides several examples of pitfalls. Dealing with definitional pitfalls has been, to date, a trial and error process where market creators learn from past mistakes and ad hoc ‘legalese’ (e.g., a ‘consensus of credible reporting’ may be used to resolve markets) is copied from market to market [1]. Future research could develop machine-checkable predicate specifications (precedence rules, ranked sources, time semantics, and default outcomes) and verify they are well-defined with model checking (cf. [15]).
@@ 3.3 Share Structure and Pricing
$R_{j_{\mathsf{k}}}(\omega)=\begin{cases}1,&\text{if }\omega=\mathsf{k},\\ 0,&\text{otherwise.}\end{cases}$ (3)
For a WTA market to be well-functioning, two conditions must hold on outcome shares. (i) They should be mutually exclusive so no more than one share wins: $R_{j_{\mathsf{k}}}(\omega)R_{j_{\mathsf{\ell}}}(\omega)=0\quad\forall\omega\in\Omega,\forall\mathsf{k}\neq\mathsf{\ell}$; and (ii) they should be complete so at least one share wins: $\sum_{\mathsf{k}\in\Omega}R_{j_{\mathsf{k}}}(\omega)=1\quad\forall\omega\in\Omega$. If they are not mutually exclusive, the operator could be under-collateralized for making all payments (in violation of solvency in Definition 3). If they are incomplete, a deficient market might end with all participants receiving \$0. A consequence is that holding one share for each outcome is equivalent to holding \$1, a fact we will return to in §3.4.
@@
The second structure we term a yes-no bundle (YNB). YNB markets were popularized by InTrade [37]. A YNB market issues two outcome shares for each outcome, a ‘yes’ and a ‘no:’ $J=\{j_{\mathsf{A_{Y}}},j_{\mathsf{A_{N}}},j_{\mathsf{B_{Y}}},j_{\mathsf{B_{N}}},j_{\mathsf{C_{Y}}},j_{\mathsf{C_{N}}}\}$. For any $k\in\{\mathsf{A},\mathsf{B},\mathsf{C}\}$,
$R_{j_{\mathsf{k_{Y}}}}(\omega)=\begin{cases}1,&\text{if }\omega=\mathsf{k},\\ 0,&\text{otherwise.}\end{cases}\qquad R_{j_{\mathsf{k_{N}}}}(\omega)=\begin{cases}1,&\text{if }\omega\neq\mathsf{k},\\ 0,&\text{otherwise.}\end{cases}$ (4)
@@
$j_{\mathsf{k_{N}}}\equiv\sum_{\ell\in\Omega\setminus\{\mathsf{k}\}}j_{\mathsf{\ell_{Y}}}\quad\text{for any }\mathsf{k}\in\Omega.$ (5)
And multiple No outcome shares can be converted into Yes shares plus cash:
$\sum_{\ell\in\Omega\setminus\{\mathsf{k}\}}j_{\mathsf{\ell_{N}}}\equiv j_{\mathsf{k_{Y}}}+\mathdollar 1\cdot(|\Omega|-2)$ (6)
@@
$R_{j_{\mathrm{lin}}}(\omega)=\begin{cases}0,&X(\omega)\leq a,\\ \dfrac{X(\omega)-a}{\,b-a\,},&a<X(\omega)<b,\\ 1,&X(\omega)\geq b.\end{cases}$ (7)

### v2

@@ Abstract
Decentralized prediction markets (DePMs) allow open participation in event-based wagering without fully relying on centralized intermediaries. We review the history of DePMs which date back to 2011 and includes hundreds of proposals. Perhaps surprising, modern DePMs like Polymarket deviate materially from earlier designs like Truthcoin and Augur v1. We use our review to present a modular workflow comprising eight stages: underlying infrastructure, market topic, share structure and pricing, initialization, trading, market resolution, settlement, and archiving. For each module, we enumerate the design variants, analyzing trade-offs around decentralization, expressiveness, and manipulation resistance. We also identify open problems for researchers interested in this ecosystem.
@@ 1 Introduction
Research Questions.
Noticing these points of differences inspired us to enumerate all of the design decisions involved in creating a DePM. What are the core modules that appear across DePM designs? How do DePMs differ in how they implement each module (creation, trading, resolution, settlement)? Which early design choices shaped later DePM architectures? Which design choices create recurring failure modes or operational problems? What trade-offs do DePMs make between decentralization, usability, and correctness of resolution? We also ask what open design gaps remain for building robust DePMs (these are marked as ‘research gaps’ in the paper).
Systemization of Knowledge (SoK).
Within the blockchain literature, a variety of SoKs have been published with a variety of methodologies and deliverables, making it difficult to establish what constitutes an SoK paper. We review several SoKs from FC and AFT that, like us, consider a blockchain technology: oracles [27], stablecoins [59], DeFi services [82], light clients [16], layer 2s [36], privacy technologies [8], and MEV mitigations [41]. Like us, every one of these papers compares/contrasts a set of technologies [27, 59, 82, 36] or theoretical models [16, 8, 41] from academia and industry. Like us, the main deliverable is decomposing designs into a taxonomy or classification system for comparison and evaluation [27, 59, 82, 36, 16, 8, 41]. Like us, some distill the technology into a single model into layers or abstractions before classifying [27], others offer definitions [16, 82]. Only one of these papers does quantitative experiments [82], a welcome but unnecessary requirement for SoKs. Unlike a literature review—which primarily summarizes prior work paper-by-paper—an SoK imposes a common set of concepts, abstractions, and evaluation criteria so that disparate systems can be compared on a level playing field rather than merely described.
Methodology.
We obtained a collection of academic works on decentralized prediction markets, as well as various intersecting topics including (centralized) prediction markets, oracles, DeFi, and AMMs. We used our knowledge of the field, Google Scholar (search and cited by features), and citations within papers. Our library is available, sorted by topic, on Zotero. [^2] Zotero: https://www.zotero.org/groups/5750510/2024_prediction_markets/library We also identified projects without white-papers or academic papers, reviewing the 97 current projects listed on a recently released online dashboard [73] and by studying 20+ historical systems that are no longer active. We also searched news sources for opinions and issues on leading decentralized prediction markets, such as Polymarket. We applied the following exclusion criteria to remove from consideration designs that are: fully permissioned CePMs, alternative wager structures beyond the first notable design we saw (e.g., parimutuel betting [78, 49, 80], back/lay exchanges [20] and cash for difference contracts [26]), supporting/adjacent infrastructure, and insufficiently documented or scams. We reviewed over a hundred projects, distilled into the 35+ notable DePMs listed in Appendix 0.A.
A modular workflow (cf. [27]) is an abstraction that breaks competing designs into the same common set of steps and parts, enabling side-by-side comparison. We used multiple rounds of affinity diagramming to cluster DePM sub-components from our corpus into eight stages: underlying infrastructure, market topic, share structure and pricing, trading, market resolution, settlement, and archiving. Relevant artifacts and data are available on GitHub. [^3] GitHub: https://github.com/MadibaGroup/2024-PredictionMarkets
@@ 2 Preliminaries
2.1 Wagering Systems
@@
A wager is a two-party contract with payouts based on the outcome of a future event. Consider Alice and Bob who wager on the same outcome of an event. With a fixed-odds bookmaker (or online betting), Alice’s contract is different from Bob’s contract in at least two regards: (i) it specifically names Alice as the counterparty and (ii) the payouts could be different if the odds changed between Alice’s wager and Bob’s. By contrast, in a prediction market contract (called a outcome share), Alice and Bob hold identical contracts: (i) all contracts are between the market operator and whoever redeems the contract, and (ii) the payout is exactly the same (typically \$0 if incorrect and \$1 if correct). Odds are reflected in the price paid for a prediction market contract (i.e., variable cost and fixed payout), while a bookmaker contract has a fixed cost and variable payout. Thus the key distinction is that prediction market outcome shares are fungible and can be freely traded between participants, enabling a free market that communicates information to the public through outcome share prices, trading volume, market depth, and other financial market metrics. Operating a prediction market also differs from operating a bookmaker in a very important way:
Insight 1. Bookmakers quote odds without needing to hold sufficient funds in the event of a payout. It is merely a promise to pay. By contrast, a prediction market is structured with sufficient collateral for all outcomes at all times (including the very first trade! See §3.4). Every dollar paid to a winning trader is a dollar from either (a) a losing trader or (b) bounded within a special loss fund setup before the market opens.
@@ 2.2 Definitions
A (single) market is a tuple $M=(\Omega,J,R)$, where $\Omega$ is a nonempty outcome space, $J$ is a finite index set of contract labels (“shares”), and $R:\Omega\to\mathbb{R}_{\geq 0}^{J}$ is a nonnegative payoff map. Intuitively, $M$ is associated with an underlying uncertain event, whose possible resolutions are represented by $\Omega$. When $M$ resolves to $\omega\in\Omega$, one unit of share $j\in J$ pays $R(\omega)_{j}$ units of the numeraire $\mathcal{N}$ (Def. 2).
@@
For a market $M=(\Omega,J,R)$, suppose there exists a bijection $\iota:J\to\Omega$ and $R(\omega)_{j}\in\{0,1\}$ with $\sum_{j\in J}R(\omega)_{j}=1$ for all $\omega\in\Omega$. Then $M$ is a winner-take-all (WTA) or Arrow–Debreu market: a unit claim of label $j$ pays $1$ iff the realized outcome equals $\iota(j)$, and 0 otherwise.
@@
A prediction–market system is a tuple $\mathcal{S}=(\mathcal{M},\mathcal{N},\mathsf{res})$, where $\mathcal{M}$ is a countable set of markets, $\mathcal{N}$ is a numeraire (unit of account), and $\mathsf{res}$ is a resolution function such that, for each $M=(\Omega,J,R)\in\mathcal{M}$, we have $\mathsf{res}(M)\in\{\bot\}\cup\Omega$, $\mathsf{res}(M)$ is initially $\bot$, and $\mathsf{res}(M)$ transitions exactly once to some $\omega\in\Omega$.
@@
For every market $M=(\Omega,J,R)\in\mathcal{M}$ operating in system $\mathcal{S}=(\mathcal{M},\mathcal{N},\mathsf{res})$, the following axioms hold. In what follows, all quantities are for this fixed market $M$, and we write $\mathsf{Treas}:=\mathsf{Treas}_{M}$ and $L:=L_{M}$.
@@
    Issuance and Solvency. Let $S=(S_{j})_{j\in J}\in\mathbb{R}_{\geq 0}^{J}$ be the outstanding supply vector, and let $\mathsf{Treas}\geq 0$ be the market’s treasury (in the numeraire $\mathcal{N}$). Define the worst-case liability
    $L(S)\;:=\;\sup_{\omega\in\Omega}\;\sum_{j\in J}S_{j}\,R(\omega)_{j}.$ (1)
    The system maintains, at all times,
    $\mathsf{Treas}\;\geq\;L(S).$ (2)
    Any operation that changes supplies (including issuance) must preserve solvency. In particular, if issuance changes supplies from $S$ to $S^{\prime}$, then treasury increases sufficiently:
    $\Delta\mathsf{Treas}\;\geq\;L(S^{\prime})-L(S).$ (3)
@@
    Transfer and Fungibility. Holdings are transferable between accounts and transfers conserve per–label totals $S_{j}$. Shares are fungible within each label: for all $j\in J$, any $\omega\in\Omega$, and any $q\geq 0$, redeeming $q$ units pays $q\,R(\omega)_{j}$ units of $\mathcal{N}$.
@@
    Settlement and redemption. If $\mathsf{res}(M)=\omega\in\Omega$, then any holder may redeem: for all $j\in J$ and $q\geq 0$, redeeming $q$ units of label $j$ pays $q\,R(\omega)_{j}$ units of $\mathcal{N}$. Redemption burns the redeemed shares and debits the treasury by the corresponding payout, preserving the solvency invariant above.
@@ 2.3 An Example of a Market
Also of interest is how the prediction market did not obviously extract insider information which is in violation of what theory would predict [39]. The director did state he did not participate in the market and advised his team working on the film not to either [10]. Friction for novice users is also high: web3 apps have a learning curve and if insiders were based at HBO in the US, access would require circumvention of Polymarket’s geofencing. Perhaps these reasons kept insiders out of the market for the 5 days it ran. By contrast, allegations of insider trading have be levelled in other markets, including the removal of Nicolas Maduro from power immediately before his seizure by the United States military [2].
Research Gap 1. Can insider trades be classified from blockchain data with reasonable percision? Are occurrences higher on unregulated markets? Are markets with insider trading more accurate? What correlations exist between insider trading activity and market metrics (price, liquidity, etc.)?
@@ 3 Modular Workflow
Table 2: Classification of selected (for significance or novelty) DePMs using our modular workflow. Keywords used in the table are defined in referenced section, while ●, ◐, ○ respectively infer that the related property is fully, partially or not supported.
We now turn to the design landscape of DePMs and step through our modular workflow (cf. [27]) of eight stages: underlying infrastructure, market topic, share structure and pricing, trading, market resolution, settlement, and archiving. For each stage, we enumerate the possible designs and discuss competing trade-offs, summarized as the columns of Table 2.
Research Gap 2. Ca accuracy metrics [33] be used to compare various CePM and DePM platforms? Are there meaningful correlations between accuracy and the market microstructure options in Table 2?
@@ 3.1 Underlying Infrastructure
Research Gap 3. On-chain markets provide open information about the trading strategies of other traders who do not deliberately and successfully obfuscate their trades and identities. Can we classify copy traders and measure occurrences? Is copy trading profitable? Under what conditions? Is there a negative impact on market accuracy due to cascading/herding effects?
@@ 3.2 Market Topic
Table 3: Some pitfalls that illustrate the difficulty in properly defining a prediction market topic, collected through our literature review and manual observation of top markets on Polymarket over 2024–2025.
@@
Insight 2. Careful attention must be paid to both the general topic of the market and the ‘fine-print’ or exact predicate that decides the market. Table 3 provides several examples of pitfalls. Dealing with definitional pitfalls has been, to date, a trial and error process where market creators learn from past mistakes and ad hoc ‘legalese’ (e.g., a ‘consensus of credible reporting’ may be used to resolve markets) is copied from market to market [1].
Research Gap 4. Can market topics be written with machine-checkable predicate specifications (precedence rules, ranked sources, time semantics, and default outcomes), where model checking (cf. [18]) could eliminate ambiguities, corner cases, and loopholes? What would real world case studies of markets with additional context tell us?
@@ 3.3 Share Structure and Pricing
$R(\omega)_{j_{\mathsf{k}}}=\begin{cases}1,&\text{if }\omega=\mathsf{k},\\ 0,&\text{otherwise.}\end{cases}$ (4)
For a WTA market to be well-functioning, two conditions must hold on outcome shares. (i) They should be mutually exclusive so no more than one share wins: $R(\omega)_{j_{\mathsf{k}}}\,R(\omega)_{j_{\mathsf{\ell}}}=0\quad\forall\omega\in\Omega,\forall\mathsf{k}\neq\mathsf{\ell}$; and (ii) they should be complete so at least one share wins: $\sum_{\mathsf{k}\in\Omega}R(\omega)_{j_{\mathsf{k}}}=1\quad\forall\omega\in\Omega$. If they are not mutually exclusive, the operator could be under-collateralized for making all payments (in violation of solvency in Definition 3). If they are incomplete, a deficient market might end with all participants receiving \$0. A consequence is that holding one share for each outcome is equivalent to holding \$1, a fact we will return to in §3.5.
@@
The second structure we term a yes-no bundle (YNB). YNB markets were popularized by InTrade [43]. A YNB market issues two share labels for each outcome, a ‘yes’ and a ‘no:’ $J=\{j_{\mathsf{A_{Y}}},j_{\mathsf{A_{N}}},j_{\mathsf{B_{Y}}},j_{\mathsf{B_{N}}},j_{\mathsf{C_{Y}}},j_{\mathsf{C_{N}}}\}$. For any $k\in\{\mathsf{A},\mathsf{B},\mathsf{C}\}$,
$R(\omega)_{j_{\mathsf{k_{Y}}}}=\begin{cases}1,&\text{if }\omega=\mathsf{k},\\ 0,&\text{otherwise.}\end{cases}\qquad R(\omega)_{j_{\mathsf{k_{N}}}}=\begin{cases}1,&\text{if }\omega\neq\mathsf{k},\\ 0,&\text{otherwise.}\end{cases}$ (5)
@@
$\sum_{\mathsf{k}\in\Omega}R(\omega)_{j_{\mathsf{k_{Y}}}}=1\quad\text{and}\quad\sum_{\mathsf{k}\in\Omega}R(\omega)_{j_{\mathsf{k_{N}}}}=|\Omega|-1\quad\text{for all }\omega\in\Omega.$ (6)
@@
$R(\omega)_{j_{\mathsf{k_{N}}}}\;=\;\sum_{\ell\in\Omega\setminus\{\mathsf{k}\}}R(\omega)_{j_{\ell_{Y}}}\quad\text{for all }\omega\in\Omega\text{ and any }\mathsf{k}\in\Omega.$ (7)
While Yes has equivalence to:
$R(\omega)_{j_{k_{Y}}}\;=\;\sum_{\ell\in\Omega\setminus\{k\}}R(\omega)_{j_{\ell_{N}}}\;-\;(|\Omega|-2)$ (8)
Research Gap 5. How does a negRisk gadget encourage price parity between equivalent positions? Can we quantify its improvements for price execution and trade execution costs? Does a reverse gadget, for converting a single yes share, offer similar benefits?
@@
$R(\omega)_{j_{\mathrm{lin}}}=\begin{cases}0,&X(\omega)\leq a,\\ \dfrac{X(\omega)-a}{\,b-a\,},&a<X(\omega)<b,\\ 1,&X(\omega)\geq b.\end{cases}$ (9)
@@
Research Gap 6. The academic literature has explored advanced market types, including combinatorial markets and establishing conditional probabilities from markets. Implementing and experimenting with these designs for DePMs is open. In sports-betting, chaining together multiple outcomes in an all-or-nothing bet (parlay) is popular but a treatment for DePMs is lacking.
3.4 Market Initialization
A near universal difference between any CePM and DePM is that a DePM allows outcome shares to be withdrawn from the platform, typically in a form compliant with a token standard such as ERC-20 or the more efficient ERC-1155 [52]. Withdrawing outcome shares allows traders to exchange tokens outside of the platform and to compose with third party DeFi services, including on-chain trading, lending, and leverage [82]. Options for trading outcome shares can be broken into two steps: (i) how does the first outcome share come into existence and how does the first trader trade, and (ii) how do traders trade once a market has been established? Probably the greatest evolution in DePMs, from Truthcoin to Polymarket, concerns how the first trade happens. There are three options: automated bookmaking, splitting, and matching.
Automated bookmaking was popularized through the academic work of Robin Hanson [38, 40], researched for Intrade [55], and first suggested for DePMs by Truthcoin [77], which heavily influenced Augur v1 [64].
Definition 4(Automated bookmaking)
Fix a market $M=(\Omega,J,R)$. An automated bookmaker posts (i) a loss fund $\mathsf{Treas}\in\mathbb{R}_{\geq 0}$ in the numeraire $\mathcal{N}$ and (ii) a pricing rule (odds) for labels $j\in J$. It accepts deposits in $\mathcal{N}$ and issues shares, updating the supply vector $S$, and must always satisfy solvency (Definition 3).
In this model, the operator sets initial prices for each outcome share (equivalent to setting market odds) and collateralizes enough payout money to cover a worst-case loss in its treasury. If Alice is the first trader, she can immediately trade with the operator. The operator is autonomous and sets buy/sell prices algorithmically, originally using Hanson’s logarithmic market scoring rule (LMSR) [40]. The key point is that the operator’s fund is Alice’s counterparty; if Alice wins, the fund loses, and vice-versa. The pros are instant liquidity for the first trader and the cons are the risk to the operator of losing money and the burden of needing to set initial odds (getting them wrong increases the chances it loses).
Acute readers might wonder the relationship between automated bookmaking and an automated market makers (e.g., Uniswap), which are discussed in the next section. Roughly speaking, a WTA market run by automated bookmaking is termed a cost-function prediction market (CFPM) and a CFPM is equivalent in pricing (and trade costs) to an AMM (defined by a set of axioms) with the right invariant [30].
Definition 5(Splitting and merging)
Fix a market $M=(\Omega,J,R)$ in numeraire $\mathcal{N}$. A splitting/merging mechanism provides counterparty-free state transitions for a fixed unit bundle $B$ of WTA share labels such that
$\sum_{j\in B}R(\omega)_{j}=1\quad\text{for all }\omega\in\Omega.$
Splitting maps $1\cdot\mathcal{N}\mapsto\sum_{j\in B}j$ (minting one unit of each $j\in B$), and merging maps $\sum_{j\in B}j\mapsto 1\cdot\mathcal{N}$ (burning them), typically before resolution. Repeating the operation yields multiple bundles.
In YNB markets, each outcome’s YES/NO pair is its own two-outcome WTA market. Splitting is per outcome: converting \$1 of collateral mints one $j_{\mathsf{k_{Y}}}$ and one $j_{\mathsf{k_{N}}}$ for the chosen $\mathsf{k}$. Across outcomes, supplies are uncoupled—the total minted for the outcome $\mathsf{A}$’s bundle need not match that for outcomes $\mathsf{B}$ or $\mathsf{C}$.
Insight 3. If the first trader makes one dollar, someone must lose a dollar. It either needs to be the operator or a second trader. This forces DePMs to choose between needing a loss fund or having no instant liquidity for the first trader.
The third approach, matching, was used by InTrade [43] and a variant by Fairlay [28], and is used today by Kalshi.
Definition 6(Matching)
Fix a market $M=(\Omega,J,R)$ in numeraire $\mathcal{N}$. For each label $j\in J$, define its per-unit worst-case payoff
$c_{j}\;:=\;\sup_{\omega\in\Omega}R(\omega)_{j},$
and define the complementary label $\bar{j}$ with payoff
$R(\omega)_{\bar{j}}\;:=\;c_{j}-R(\omega)_{j}.$
A matching mechanism accepts orders of the form $(j,q,p)$, meaning “trade $q\geq 0$ units of label $j$ at price $p$ (in $\mathcal{N}$) per unit,” and executes a trade only when a buy order and a sell order coincide on $(j,q,p)$. Upon such a match, the mechanism: (i) transfers $qp$ units of $\mathcal{N}$ from the buyer and $q(c_{j}-p)$ units of $\mathcal{N}$ from the seller into escrow; (ii) mints $q$ units of label $j$ to the buyer and $q$ units of label $\bar{j}$ to the seller. At resolution $\omega\in\Omega$, the two positions pay $qR(\omega)_{j}$ and $qR(\omega)_{\bar{j}}$, which sum to $qc_{j}$, exactly covered by the escrow; hence the operator is riskless.
Briefly, it mirrors a futures market, where Alice posts a desired short/long position at a chosen price on an orderbook with a margin account holding enough cash to cover her maximum loss if she obtains the position. If Bob is willing to take the other side, also with sufficient margin for his maximum loss, the operator matches them, creates two shares and gives them to Alice and Bob. Alice and Bob are not counterparties, both settle with the operator once the shares are created, however their coincidence of wants (COW) is necessary for the operator to create shares at no risk to itself.
Insight 4. A DePM operated through smart contracts should use automated bookmaking or splitting/merging. Matching effectively requires a central limit order book, which is too expensive to run on-chain at scale [60].